# Beyond Additivity: Sparse Isotonic Shapley Regression toward Nonlinear Explainability

### v1

@@ Abstract
Shapley values, a gold standard for feature attribution in Explainable AI, face two primary challenges. First, the canonical Shapley framework assumes that the worth function is additive, yet real-world payoff constructions—driven by non-Gaussian distributions, heavy tails, feature dependence, or domain-specific loss scales—often violate this assumption, leading to distorted attributions. Secondly, achieving sparse explanations in high-dimensional settings by computing dense Shapley values and then applying ad hoc thresholding is prohibitively costly and risks inconsistency. We introduce Sparse Isotonic Shapley Regression (SISR), a unified nonlinear explanation framework. SISR simultaneously learns a monotonic transformation to restore additivity—obviating the need for a closed-form specification—and enforces an L0 sparsity constraint on the Shapley vector, enhancing computational efficiency in large feature spaces. Its optimization algorithm leverages Pool-Adjacent-Violators for efficient isotonic regression and normalized hard-thresholding for support selection, yielding ease in implementation and global convergence guarantees. Analysis shows that SISR recovers the true transformation in a wide range of scenarios and achieves strong support recovery even in high noise. Moreover, we are the first to demonstrate that irrelevant features and inter-feature dependencies can induce a true payoff transformation that deviates substantially from linearity. Extensive experiments in regression, logistic regression, and tree ensembles demonstrate that SISR stabilizes attributions across payoff schemes, correctly filters irrelevant features; in contrast, standard Shapley values suffer severe rank and sign distortions. By unifying nonlinear transformation estimation with sparsity pursuit, SISR advances the frontier of nonlinear explainability, providing a theoretically grounded and practical attribution framework.
@@ 1 Introduction and Motivation
$V_{A}(\{\beta_{j}\}_{j\in A})=\max_{j\in A}\beta_{j},\ \forall A\subseteq 2^{F}.$ (3)
@@
  • The optimization algorithm developed for SISR features simple, closed-form updates and is accompanied by global convergence guarantees. The incorporation of sparsity improves computational efficiency
@@ 2 Proposed Method
$\begin{split}\min_{\beta\in\mathbb{R}^{p},c\in\mathbb{R}}&\sum_{A\subseteq 2^{F},\,A\neq\emptyset,\,A\neq F}w_{\text{\tiny SH}}(A)\left(\nu_{A}-\sum_{j\in A}\beta_{j}-c\right)^{2}\\ &\text{subject to}\quad c=\nu_{\emptyset},\quad c+\sum_{j=1}^{p}\beta_{j}=\nu_{F},\end{split}$ (4)
@@
$\min_{\beta,c}$ $\sum_{A\subseteq 2^{F}}w_{\text{\tiny SH}}(A)(\nu_{A}-\sum_{j\in A}\beta_{j}-c)^{2}$
@@
$\nu_{A}^{c}=\nu_{A}-\nu_{\emptyset},\quad\forall A\subseteq 2^{F},$ (7)
@@
$\min_{\beta}\sum_{A\subseteq 2^{F}}w_{\text{\tiny SH}}(A)\left(\nu_{A}^{c}-\sum_{j\in A}\beta_{j}\right)^{2}.$
@@
In our view, one viable solution is to apply a transformation that promote Gaussianity. Let’s consider an alternative Shapley value model in a transformed domain:
@@
$\min_{\beta,T(\cdot)}\sum_{A\subseteq 2^{F}}w_{\text{\tiny SH}}(A)\Big\{T(\nu_{A})-\sum_{j\in A}T(\beta_{j})\Big\}^{2}\ \text{subject to}\ \beta\in\mathcal{C},\,T(\cdot)\in\mathcal{T},$ (11)
where the objective minimizes the Shapley-weighted sum of squared differences between the transformed coalition values $T(\nu_{A})$ and the transformed linear sum $\sum_{j\in A}T(\beta_{j})$ over all subsets $A\subseteq 2^{F}$. Here, we use $\mathcal{C}\subseteq\mathbb{R}^{p}$ to denote the constraint set for $\beta$, and $\mathcal{T}$ to denote the class of admissible transformation functions. By the notational convention for $A=\emptyset$, (11) automatically enforces
@@
In this paper, we focus a specific instance of (11), referred to as the “Sparse Isotonic Shapley Regression” (SISR):
$\begin{split}&\min_{\beta,T(\cdot)}\sum_{A\subseteq 2^{F}}w_{\text{\tiny SH}}(A)\Big\{T(\nu_{A})-\sum_{j\in A}T(\beta_{j})\Big\}^{2}\\ &\mbox{ subject to }\|\beta\|_{0}\leq s,T\in\mathcal{M},\sum_{j=1}^{p}(T(\beta_{j}))^{2}=1,\end{split}$ (13)
@@
A normalization is imposed on the transformed feature contributions, $\sum_{j=1}^{p}(T(\beta_{j}))^{2}=1.$ This prevents degeneracy (e.g., trivial solutions such as $T\equiv 0$) and anchors the scale of the model. An appealing feature of (13) is its invariance to the overall scaling of $\{\nu_{A}\}$, and the normalization constant is fixed at 1 without loss of generality. Moreover, Section 3 will show that imposing such a spherical constraint yields computational benefits, enabling a closed-form solution for the attribution-update and improving implementability.
@@
$\min_{\gamma,T(\cdot)}\quad\sum_{A\subseteq 2^{F}}w_{\text{\tiny SH}}(A)(T(\nu_{A})-\sum_{j\in A}\gamma_{j})^{2}\text{ s.t. }\|\gamma\|_{0}\leq s,\|\gamma\|_{2}=1,T\in\mathcal{M}.$ (15)
The corresponding model assumption is thus $T^{*}(\nu_{A})\sim\mathcal{N}(\sum_{j\in A}\gamma_{j}^{*},\ \sigma_{A}^{2})$ for all $A\subseteq 2^{F},$ where the genuine transformation function $T^{*}$ is monotonic with $T^{*}(0)=0$, and $\nu_{A}$ are assumed to be independent across different subsets $A$. The starred quantities represent the underlying statistical truth of interest to estimate. Assume $\gamma^{*}\in\mathbb{R}^{p}$ satisfies $\|\gamma^{*}\|_{0}\leq s^{*}$ and $\|\gamma^{*}\|_{2}=1$ with $1\leq s^{*}\leq p$ and $s$ is specified as an upper bound on $s^{*}$. After estimating $\hat{\gamma}$ and $\hat{T}$ from (15), one can recover the $\beta$-scores by applying the inverse transformation $\hat{\beta}_{j}=\hat{T}^{-1}(\hat{\gamma}_{j}).$ This reconstructs the multivariate relationship between $\nu_{A}$ and the set of feature contributions in the original scale, yielding $\nu_{A}\approx\hat{T}^{-1}(\sum_{j\in A}\hat{T}(\hat{\beta}_{j}))$, to offer interpretable Shapley-based attributions.
@@ 3 Optimization Algorithm
$\nu=\big[\nu_{A}\big]_{A\subseteq 2^{F}}\in\mathbb{R}^{2^{p}},\quad\delta=\Big[\sum_{j\in A}\gamma_{j}\Big]_{A\subseteq 2^{F}}=Z\gamma\in\mathbb{R}^{2^{p}},$
where $Z\in\mathbb{R}^{2^{p}\times p}$ is the “incidence matrix” indicating which features are active in each subset $A$, aligned with the same ordering used for $t$. Henceforth, we also write $\nu_{i}$ (and likewise $\delta_{i}$) to denote the entry corresponding to the $i$th subset. Additionally, introduce the diagonal weight matrix
$W=\text{diag}\{w_{\text{\tiny SH}}(A)\}_{A\subseteq 2^{F}}.$ (16)
@@
where $E$ encodes the pairwise ordering constraints induced by $\nu$, due to the monotonicity of the transformation $T$. This formulation replaces strict monotonicity with a non-decreasing constraint, a mild adjustment that facilitates numerical implementation. In the following of the section, we design a two-block alternating optimization algorithm.
@@
Let $\rho\geq\|Z^{\top}WZ\|_{2}$, where $\|\cdot\|_{2}$ denotes the martrix spectral norm. For any initial point $\gamma^{(0)}$ satisfying $\|\gamma^{(0)}\|_{0}\leq s$ and $\|\gamma^{(0)}\|_{2}=1$, the sequence $\{\gamma^{(k)}\}$ generated by (21) produces non-increasing (and thus convergent) function values:
@@
Input: $\nu=[\nu_{A}]_{A\subseteq 2^{F}}\in\mathbb{R}^{2^{p}}$ (baseline-adjusted, such that $\nu_{\emptyset}=0$), sparsity level $s$, design matrix $Z\in\mathbb{R}^{2^{p}\times p}$, diagonal weight matrix $W$ (cf. (16)), and an initial vector $t^{(0)}\in\mathbb{R}^{2^{p}}$ (e.g., $C\nu$ with a large $C$ if $\|\nu\|_{\infty}$ is small, to improve precision, and $C=1$ otherwise).
@@ 4.1 Domain Adaptation
Figure 1 presents the results under 6 different functional forms for the true transformation $T^{*}$: square root ($T^{*}=(\cdot)^{1/2}$), fifth root ($T^{*}=(\cdot)^{1/5}$), exponential ($T^{*}=\exp(\cdot)-1$), logarithmic ($T^{*}=\log(\cdot+1)$), tangent ($T^{*}=\tan(\cdot)/c_{1},c_{1}=10$), and normal distribution ($T^{*}=\Phi(\cdot+c_{2})/c_{1},Q(\cdot)=\Phi^{-1}(c_{1}\cdot)-c_{2},c_{1}=1/\sqrt{3},c_{2}=Q(c_{1}\sigma_{\min})$). As pointed out by a reviewer, comparing the estimated $\hat{T}$ directly to the true $T^{*}$, rather than their inverses on the $\nu$ scale, is statistically sound, as this aligns with our additive Gaussian assumption (9). Encouragingly, across all cases, the estimated transformation $\hat{T}(\nu)$ closely aligns the ground-truth $T^{*}(\nu)$, providing strong empirical evidence for the effectiveness of SISR in accurately recovering the underlying transformation structure.
@@ 4.7 Diabetes
Using the negative-entropy payoff, the estimated transformation is piecewise linear—an indicator of sparsity in the underlying feature importances (cf. Section 4.3). Although the segment slopes differ only modestly, the resulting shift alters several attributions: the largest discrepancy is for DiabetesPedigreeFunction, whose importance is noticeably lower under the standard Shapley calculation than in the SISR-calibrated version. The bottom panel reports results for the likelihood payoff. Here, the conventional Shapley values are severely distorted: numerous features appear almost as influential as Glucose, and Pregnancies even becomes negative, illustrating how conventional attributions can be distorted by such an utility function. By contrast, the SISR estimates without sparsity nearly replicate those obtained under negative entropy, and introducing sparsity leaves the ranking largely unchanged. Overall, SISR delivers stable attributions across different value (payoff) functions, whereas the standard procedure is highly sensitive to which value function is chosen.
@@ 5 Discussion and Extensions
$\sum_{A\subseteq 2^{F}}w_{\text{\tiny SH}}(A)L(\delta_{A};T(\nu_{A})),\quad\delta_{A}=\sum_{j\in A}T(\beta_{j}),\quad L(\eta,y)=-\eta y+b(\eta),$
where $b$ is the log-partition function in the specified natural exponential distribution and $b^{-1}$ induces the canonical link (Hardin and Hilbe,, 2018). The resulting optimization subject to the same monotonicity constraints and sparsity/normalization constraints as in (13) may be termed as the sparse isotonic Shapley GLM. When $b(\eta)=\eta^{2}/2$, this reduces to the squared-loss formulation analyzed in previous sections.
An appealing fact is that the overall computational framework in Section 3 carries over seamlessly. The update of $t$ (and hence $\delta$) is obtained by an order-restricted fit using the isotonic machinery adapted to the GLM loss; see, e.g., de Leeuw et al., (2009) and Luss and Rosset, (2014) for related PAVA and active-set extensions. The update of $\gamma$ uses exactly the same surrogate function as in (20) with the working loss $l(\gamma)=\sum_{A\subseteq 2^{F}}w_{\text{\tiny SH}}(A)L(\delta_{A};T(\nu_{A}))$, for which a direct calculation yields $\nabla l(\gamma)=Z^{\top}W(b^{\prime}(Z\gamma)-t)$, recovering (19) in the Gaussian case. The nonlinear operators from Theorem 1 are unchanged and Theorem 2 continues to apply with a distribution-specific curvature bound. For example, in logistic regression where $b(\eta)=\log(1+\exp(\eta))$, the lower bound on $\rho$ becomes $\|Z^{\top}WZ\|_{2}/4$ (since $b^{\prime\prime}\leq 1/4$). Derivations follow Section 3 closely and are omitted.

### v2

@@ Abstract
Shapley values, a gold standard for feature attribution in Explainable AI, face two primary challenges. First, the canonical Shapley framework assumes that the worth function is additive, yet real-world payoff constructions—driven by non-Gaussian distributions, heavy tails, feature dependence, or domain-specific loss scales—often violate this assumption, leading to distorted attributions. Second, achieving sparse explanations in high-dimensional settings by computing dense Shapley values and then applying ad hoc thresholding is prohibitively costly and risks inconsistency. We introduce Sparse Isotonic Shapley Regression (SISR), a unified nonlinear explanation framework. SISR simultaneously learns a monotonic transformation to restore additivity—obviating the need for a closed-form specification—and enforces an L0 sparsity constraint on the Shapley vector, enhancing computational efficiency in large feature spaces. Its optimization algorithm leverages Pool-Adjacent-Violators for efficient isotonic regression and normalized hard-thresholding for support selection, ensuring ease in implementation and global convergence guarantees. Analysis shows that SISR recovers the true transformation in a wide range of scenarios and achieves strong support recovery even in high noise. Moreover, we are the first to demonstrate that irrelevant features and inter-feature dependencies can induce a true payoff transformation that deviates substantially from linearity. Extensive experiments in regression, logistic regression, and tree ensembles demonstrate that SISR stabilizes attributions across payoff schemes and correctly filters irrelevant features; in contrast, standard Shapley values suffer severe rank and sign distortions. By unifying nonlinear transformation estimation with sparsity pursuit, SISR advances the frontier of nonlinear explainability, providing a theoretically grounded and practical attribution framework.
@@ 1 Introduction and Motivation
$V_{A}(\{\beta_{j}\}_{j\in A})=\max_{j\in A}\beta_{j},\ \forall A\subseteq F.$ (3)
@@
  • The optimization algorithm developed for SISR features simple, closed-form updates and is accompanied by global convergence guarantees. The incorporation of sparsity improves computational efficiency.
@@ 2 Proposed Method
$\begin{split}\min_{\beta\in\mathbb{R}^{p},c\in\mathbb{R}}&\sum_{A\in 2^{F},\,A\neq\emptyset,\,A\neq F}w_{\text{\tiny SH}}(A)\left(\nu_{A}-\sum_{j\in A}\beta_{j}-c\right)^{2}\\ &\text{subject to}\quad c=\nu_{\emptyset},\quad c+\sum_{j=1}^{p}\beta_{j}=\nu_{F},\end{split}$ (4)
@@
$\min_{\beta,c}$ $\sum_{A\in 2^{F}}w_{\text{\tiny SH}}(A)(\nu_{A}-\sum_{j\in A}\beta_{j}-c)^{2}$
@@
$\nu_{A}^{c}=\nu_{A}-\nu_{\emptyset},\quad\forall A\in 2^{F},$ (7)
@@
$\min_{\beta}\sum_{A\in 2^{F}}w_{\text{\tiny SH}}(A)\left(\nu_{A}^{c}-\sum_{j\in A}\beta_{j}\right)^{2}.$
@@
In our view, one viable solution is to apply a transformation that promotes Gaussianity. Let’s consider an alternative Shapley value model in a transformed domain:
@@
$\min_{\beta,T(\cdot)}\sum_{A\in 2^{F}}w_{\text{\tiny SH}}(A)\Big\{T(\nu_{A})-\sum_{j\in A}T(\beta_{j})\Big\}^{2}\ \text{subject to}\ \beta\in\mathcal{C},\,T(\cdot)\in\mathcal{T},$ (11)
where the objective minimizes the Shapley-weighted sum of squared differences between the transformed coalition values $T(\nu_{A})$ and the transformed linear sum $\sum_{j\in A}T(\beta_{j})$ over all subsets $A\in 2^{F}$. Here, we use $\mathcal{C}\subseteq\mathbb{R}^{p}$ to denote the constraint set for $\beta$, and $\mathcal{T}$ to denote the class of admissible transformation functions. By the notational convention for $A=\emptyset$, (11) automatically enforces
@@
In this paper, we focus on a specific instance of (11), referred to as the “Sparse Isotonic Shapley Regression” (SISR):
$\begin{split}&\min_{\beta,T(\cdot)}\sum_{A\in 2^{F}}w_{\text{\tiny SH}}(A)\Big\{T(\nu_{A})-\sum_{j\in A}T(\beta_{j})\Big\}^{2}\\ &\mbox{ subject to }\|\beta\|_{0}\leq s,T\in\mathcal{M},\sum_{j=1}^{p}(T(\beta_{j}))^{2}=1,\end{split}$ (13)
@@
A normalization is imposed on the transformed feature contributions, $\sum_{j=1}^{p}(T(\beta_{j}))^{2}=1.$ This prevents degeneracy (e.g., trivial solutions such as $T\equiv 0$) and anchors the scale of the model. An appealing feature of (13) is its invariance to the overall scaling of $\{\nu_{A}\}$, and the normalization constant is fixed at 1 without loss of generality. Moreover, Section 3 will show that imposing such a spherical constraint yields computational benefits, enabling a closed-form solution for the attribution update and improving implementability.
@@
$\min_{\gamma,T(\cdot)}\quad\sum_{A\in 2^{F}}w_{\text{\tiny SH}}(A)(T(\nu_{A})-\sum_{j\in A}\gamma_{j})^{2}\text{ s.t. }\|\gamma\|_{0}\leq s,\|\gamma\|_{2}=1,T\in\mathcal{M}.$ (15)
The corresponding model assumption is thus $T^{*}(\nu_{A})\sim\mathcal{N}(\sum_{j\in A}\gamma_{j}^{*},\ \sigma_{A}^{2})$ for all $A\in 2^{F},$ where the genuine transformation function $T^{*}$ is monotonic with $T^{*}(0)=0$, and $\nu_{A}$ are assumed to be independent across different subsets $A$. The starred quantities represent the underlying statistical truth of interest to estimate. Assume $\gamma^{*}\in\mathbb{R}^{p}$ satisfies $\|\gamma^{*}\|_{0}\leq s^{*}$ and $\|\gamma^{*}\|_{2}=1$ with $1\leq s^{*}\leq p$ and $s$ is specified as an upper bound on $s^{*}$. After estimating $\hat{\gamma}$ and $\hat{T}$ from (15), one can recover the $\beta$-scores by applying the inverse transformation $\hat{\beta}_{j}=\hat{T}^{-1}(\hat{\gamma}_{j}).$ This reconstructs the multivariate relationship between $\nu_{A}$ and the set of feature contributions in the original scale, yielding $\nu_{A}\approx\hat{T}^{-1}(\sum_{j\in A}\hat{T}(\hat{\beta}_{j}))$, to offer interpretable Shapley-based attributions.
@@ 3 Optimization Algorithm
$\nu=\big[\nu_{A}\big]_{A\in 2^{F}}\in\mathbb{R}^{2^{p}},\quad\delta=\Big[\sum_{j\in A}\gamma_{j}\Big]_{A\in 2^{F}}=Z\gamma\in\mathbb{R}^{2^{p}},$
where $Z\in\mathbb{R}^{2^{p}\times p}$ is the “incidence matrix” indicating which features are active in each subset $A$, aligned with the same ordering used for $t$. Henceforth, we also write $\nu_{i}$ (and likewise $\delta_{i}$) to denote the entry corresponding to the $i$th subset. Additionally, we introduce the diagonal weight matrix
$W=\text{diag}\{w_{\text{\tiny SH}}(A)\}_{A\in 2^{F}}.$ (16)
@@
where $E$ encodes the pairwise ordering constraints induced by $\nu$, due to the monotonicity of the transformation $T$. This formulation replaces strict monotonicity with a non-decreasing constraint, a mild adjustment that facilitates numerical implementation. In the remainder of the section, we design a two-block alternating optimization algorithm.
@@
Let $\rho\geq\|Z^{\top}WZ\|_{2}$, where $\|\cdot\|_{2}$ denotes the matrix spectral norm. For any initial point $\gamma^{(0)}$ satisfying $\|\gamma^{(0)}\|_{0}\leq s$ and $\|\gamma^{(0)}\|_{2}=1$, the sequence $\{\gamma^{(k)}\}$ generated by (21) produces non-increasing (and thus convergent) function values:
@@
Input: $\nu=[\nu_{A}]_{A\in 2^{F}}\in\mathbb{R}^{2^{p}}$ (baseline-adjusted, such that $\nu_{\emptyset}=0$), sparsity level $s$, design matrix $Z\in\mathbb{R}^{2^{p}\times p}$, diagonal weight matrix $W$ (cf. (16)), and an initial vector $t^{(0)}\in\mathbb{R}^{2^{p}}$ (e.g., $C\nu$ with a large $C$ if $\|\nu\|_{\infty}$ is small, to improve precision, and $C=1$ otherwise).
@@ 4.1 Domain Adaptation
Figure 1 presents the results under 6 different functional forms for the true transformation $T^{*}$: square root ($T^{*}=(\cdot)^{1/2}$), fifth root ($T^{*}=(\cdot)^{1/5}$), exponential ($T^{*}=\exp(\cdot)-1$), logarithmic ($T^{*}=\log(\cdot+1)$), tangent ($T^{*}=\tan(\cdot)/c_{1},c_{1}=10$), and normal distribution ($T^{*}=\Phi(\cdot+c_{2})/c_{1},Q(\cdot)=\Phi^{-1}(c_{1}\cdot)-c_{2},c_{1}=1/\sqrt{3},c_{2}=Q(c_{1}\sigma_{\min})$). As pointed out by a reviewer, comparing the estimated $\hat{T}$ directly to the true $T^{*}$, rather than their inverses on the $\nu$ scale, is statistically sound, as this aligns with our additive Gaussian assumption (9). Encouragingly, across all cases, the estimated transformation $\hat{T}(\nu)$ closely aligns with the ground-truth $T^{*}(\nu)$, providing strong empirical evidence for the effectiveness of SISR in accurately recovering the underlying transformation structure.
@@ 4.7 Diabetes
Using the negative-entropy payoff, the estimated transformation is piecewise linear—an indicator of sparsity in the underlying feature importances (cf. Section 4.3). Although the segment slopes differ only modestly, the resulting shift alters several attributions: the largest discrepancy is for DiabetesPedigreeFunction, whose importance is noticeably lower under the standard Shapley calculation than in the SISR-calibrated version. The bottom panel reports results for the likelihood payoff. Here, the conventional Shapley values are severely distorted: numerous features appear almost as influential as Glucose, and Pregnancies even becomes negative, illustrating how conventional attributions can be distorted by such a utility function. By contrast, the SISR estimates without sparsity nearly replicate those obtained under negative entropy, and introducing sparsity leaves the ranking largely unchanged. Overall, SISR delivers stable attributions across different value (payoff) functions, whereas the standard procedure is highly sensitive to which value function is chosen.
@@ 5 Discussion and Extensions
$\sum_{A\in 2^{F}}w_{\text{\tiny SH}}(A)L(\delta_{A};T(\nu_{A})),\quad\delta_{A}=\sum_{j\in A}T(\beta_{j}),\quad L(\eta,y)=-\eta y+b(\eta),$
where $b$ is the log-partition function in the specified natural exponential distribution and $(b^{\prime})^{-1}$ induces the canonical link (Hardin and Hilbe,, 2018). The resulting optimization subject to the same monotonicity constraints and sparsity/normalization constraints as in (13) may be termed as the sparse isotonic Shapley GLM. When $b(\eta)=\eta^{2}/2$, this reduces to the squared-loss formulation analyzed in previous sections.
An appealing fact is that the overall computational framework in Section 3 carries over seamlessly. The update of $t$ (and hence $\delta$) is obtained by an order-restricted fit using the isotonic machinery adapted to the GLM loss; see, e.g., de Leeuw et al., (2009) and Luss and Rosset, (2014) for related PAVA and active-set extensions. The update of $\gamma$ uses exactly the same surrogate function as in (20) with the working loss $l(\gamma)=\sum_{A\in 2^{F}}w_{\text{\tiny SH}}(A)L(\delta_{A};T(\nu_{A}))$, for which a direct calculation yields $\nabla l(\gamma)=Z^{\top}W(b^{\prime}(Z\gamma)-t)$, recovering (19) in the Gaussian case. The nonlinear operators from Theorem 1 are unchanged and Theorem 2 continues to apply with a distribution-specific curvature bound. For example, in logistic regression where $b(\eta)=\log(1+\exp(\eta))$, the lower bound on $\rho$ becomes $\|Z^{\top}WZ\|_{2}/4$ (since $b^{\prime\prime}\leq 1/4$). Derivations follow Section 3 closely and are omitted.